At 2:00pm a runner heads north on a highway, jogging at 10mph. At 2:30pm a driver heads north on the same highway to pick up the runner. If the car travels at 55mph, how long will it take the driver to catch the runner?
Answer by lwsshak3(11628)

At 2:00pm a runner heads north on a highway, jogging at 10mph. At 2:30pm a driver heads north on the same highway to pick up the runner. If the car travels at 55mph, how long will it take the driver to catch the runner?
..
Both runner and driver covered the same distance.
Let x=hours driver will catch up with runner.
Then 55x=distance driver covered
Runner ran for 1/2 hour plus the catch up time.
Then 10((1/2)+x)=distance runner covered.
10((1/2)+x)=55x
5+10x=55x
45x=5
x=1/9 hr
ans:
I would take the driver 1/9 hour to catch up with the runner.